OpenStudy (anonymous):

If 5 is subtracted from three times a number the result is 24. that is >>> 1.66?

OpenStudy (anonymous):

im sorry the result is 10 not 24

OpenStudy (anonymous):

so then its 5-3x=10 is 1.66

OpenStudy (dtan5457):

No, x would be 9.66666666667. It says 5 is subtracted FROM the number. 3x-5=24 x= 9.66666666667. Or 9 and 2/3.

OpenStudy (dtan5457):

I'm sorry, if the solution is 10. It would be 3x-5=10 x=5.

OpenStudy (anonymous):

lol oh okay i need to read carefully thank you! @dtan5457

OpenStudy (df001):

If 5 is SUBTRACTED from three times a number the result is 10. Since we don't know "a number" we use x to replace it. The equation for this is 3(x)-5=10

OpenStudy (anonymous):

thanks I just had it backwards :) @DF001

OpenStudy (dtan5457):

Make sure you didn't right 9 and 2/3. It's 5, I didn't see it equaled to 10 until after I posted.

OpenStudy (anonymous):

oh okay I will check it by the way I have to :/ thank you though! really appreciate it! @dtan5457